OpenStudy (anonymous):

multiply:: (2sqrt3+5)(3sqrt3-7)

OpenStudy (anonymous):

So you distribute as you would before (a+b)(c+d) So when you distribute 2sqrt3 to the (3sqrt3-7) what do you have? 2sqrt3(3sqrt3-7)= 2sqrt3 * 3sqrt3 -2sqrt3 * -7=6*3 - 14sqrt3=18 -14 sqrt3 Now distribute 5(3sqrt3 - 7)= 15 sqrt 3-35 You add the results 18-14 sqrt3 + 15 sqrt3 -35=-31+(-14+15)sqrt3=-31+sqrt3

OpenStudy (anonymous):

do you understand? recall that \[\sqrt3*\sqrt3=3\]

OpenStudy (anonymous):

i understand...but the answer choices are: -35-34sqrt3 -35+6sqrt3 -17+sqrt3 -17+29sqrt3

OpenStudy (anonymous):

I caught my mistake "You add the results 18-14 sqrt3 + 15 sqrt3 -35" after this it should be =18-35+(-14+15)sqrt3=-17+sqrt3

OpenStudy (anonymous):

there we go thanks youre the best!

OpenStudy (anonymous):

you are welcome!